# Toward Robust, Reproducible, and Widely Accessible Intracranial Language Brain-Computer Interfaces: A Comprehensive Review of Neural Mechanisms, Hardware, Algorithms, Evaluation, Clinical Pathways and Future Directions

### 03 Mar 2026 v1

@@ Abstract
Intracranial language brain-computer interfaces (BCIs) are a promising route for restoring communication in people with severe motor and speech impairments, but clinical translation remains limited by fragmented evidence and unresolved design trade-offs across neuroscience, hardware, algorithm, evaluation, and clinical deployment. This review synthesizes progress in neural mechanisms of overt, mimed, and imagined speech; decision-oriented hardware comparisons of microelectrode array (MEA), electrocorticography (ECoG), and stereotactic electroencephalography (SEEG) recording modalities; experiment design for cross-subject and multilingual generalization; and neural decoding advances spanning sequence models, transformers, articulatory intermediates, and language-prior-assisted frameworks. We highlight persistent bottlenecks, including weak cross-subject transfer, long-term non-stationarity and recalibration burden, heterogeneous and non-comparable evaluation practices, limited naturalistic expressivity (especially for tonal/logosyllabic languages), and low signal-to-noise ratio (SNR) of neural activity in covert speech decoding. Our contributions are threefold: (1) an end-to-end, decision-oriented synthesis linking neural representations to recording choices, experimental design, decoding model architectures, and translational constraints; (2) a structured framework organized around five coupled design questions, together with a unified evaluation framework and a cross-language/cross-task benchmark template integrating objective, perceptual, expressive, conversational, and longitudinal metrics; and (3) user-centered translational guidance covering agency-preserving shared control, verifiable performance priorities, and scenario-specific minimum viable system (MVP) profiles for reliability-first home communication versus fidelity-first conversational speech restoration. We conclude with a call for larger multilingual multi-center longitudinal datasets, harmonized benchmarks, adaptive yet interpretable decoders, prospective clinical validation, and transparent data-sharing/reporting practices under ethical protection to accelerate safe and equitable deployment of speech neuroprostheses.
@@ 1 Introduction
Speech is the most efficient and natural channel for human communication, yet it is among the first abilities to be lost in a range of devastating neurological conditions [128, 17, 85]. Individuals with locked-in syndrome (LIS), advanced amyotrophic lateral sclerosis (ALS), brainstem stroke, or severe paralysis may retain intact cognition while being unable to produce intelligible speech or reliable motor output [97, 118, 85, 23]. For these patients, language BCIs, systems that infer intended linguistic content (and ideally expressive prosody) directly from neural activity, offer a pathway to restoring communication with greater speed, privacy, and user agency than conventional assistive technologies [128, 79, 109, 80, 12].
Over the past decade, intracranial neural recordings have catalyzed rapid progress in language BCI research. Compared with non-invasive modalities [27, 140, 77, 59, 73, 44], intracranial approaches provide higher SNR and access to high-frequency activity that tightly tracks local cortical computations, enabling decoding at the levels of phonemes, words, and continuous speech [70, 69, 128]. In parallel, the field has broadened from proof-of-concept demonstrations toward system-level designs that must operate reliably in real-world settings: across days and months, across electrode configurations, across languages and speaking styles, and across the spectrum from overt speech to silently mouthed or imagined speech [78, 4]. As summarized in the lane-based timeline in Fig. 1, this evolution is not only a chronology of improved decoders, but also a shift toward chronic/online systems and frameworks that explicitly address cross-subject transfer, electrode heterogeneity, and deployment constraints. These requirements are more stringent than those for laboratory demonstrations, and they demand an explicit integration of neuroscience, hardware, machine learning, evaluation methodology, and clinical translation.
@@
On the technology side, intracranial language BCIs sit at the intersection of multiple recording modalities, each with fundamentally different trade-offs. Intracortical MEAs can capture high-dimensional spiking activity with fine spatial granularity but typically cover limited cortical territory and face chronic stability challenges [128, 12]. ECoG grids and depth electrodes such as SEEG offer broader coverage and clinically established implantation pathways, yet they sample neural activity at different spatial scales and with different noise characteristics [70, 69, 19]. The “best” modality therefore depends on the intended use case (research vs. home use), surgical feasibility, target neural features, and acceptable calibration burden. Experiment design is another critical consideration. The field has evolved from small, single-subject datasets with limited linguistic diversity toward larger, multi-subject datasets that enable cross-subject learning and benchmarking. However, challenges remain in standardizing data collection protocols, ensuring multilingual coverage, and providing rich annotations that capture not only lexical content but also prosody, emphasis, and affective features [121, 99, 34, 72, 122, 58].
@@
Algorithmically, language BCIs have evolved from feature-engineered pipelines and frame-wise classifiers to sequence-to-sequence models that jointly learn neural feature extraction, alignment, and decoding [80]. Advances such as transformers, convolutional temporal models, and hybrid architectures have enabled larger vocabularies, streaming/online decoding, and improved exploitation of long-range context. Additionally, biologically grounded articulatory intermediates provide structural constraints that enhance data efficiency and generalization [19, 18]. More recently, language models and other linguistic priors have been integrated to reduce error rates and increase output coherence, raising new questions about controllability, bias, latency, and attribution of “who said what” in shared-control system [128, 12]. These trends underscore the importance of evaluating models not only in terms of accuracy but also considering calibration requirements, real-time constraints, robustness to neural non-stationarity, and interpretability.
@@
Despite these advances, several core bottlenecks continue to hinder clinical translation of intracranial language BCIs. First, poor cross-subject generalization persists: many decoders remain subject-specific, do not transfer across electrode layouts, and cannot fully leverage data across participants or centers [19, 18, 80]. Second, long-term stability of both neural signals and end-to-end performance remains a central challenge, with MEAs often requiring frequent recalibration and ECoG/SEEG systems still needing rigorous chronic validation under naturalistic use [78, 69, 70, 4, 12]. Third, evaluation practices remain heterogeneous: inconsistent use of word error rate (WER), phoneme error rate (PER), latency, information-transfer measures, and human perceptual scoring makes cross-task comparisons difficult and can obscure clinically meaningful differences [30, 131, 20, 114]. Fourth, limited naturalistic expressivity restricts many systems to lexical content, while paralinguistic features such as prosody, intonation, emphasis, and affect are less consistently modeled; this challenge is amplified for tonal and logosyllabic languages (e.g., Mandarin, Cantonese) where pitch and syllable structure are integral to meaning [5, 99, 34, 72, 122, 58, 11, 21]. Finally, covert decoding (silent articulation or imagined speech) poses unique difficulties because neural activations tend to be weaker, more variable, and less time-locked to measurable acoustic ground truth [74, 50].
@@
Despite rapid progress, the intracranial language BCI literature remains fragmented across neuroscience, hardware, algorithm, evaluation, and clinical translation, making it difficult to compare studies, identify clinically meaningful trade-offs, and derive actionable design principles. In particular, unresolved cross-subject generalization, limited long-term stability evidence, heterogeneous evaluation metrics, and insufficient consideration of naturalistic expressivity and covert speech continue to impede translation. To address this gap, this review makes three core contributions. First, we provide an end-to-end, decision-oriented synthesis that links neural mechanisms, recording choices, experiment design, decoding models, evaluation, and translational deployment into a single framework. Fig. 2 previews this end-to-end pipeline, from neural recordings (e.g., ECoG/EEG/MEA) through feature extraction and neural decoding to outputs such as phonemes, text, speech features, and synthesized speech, while making explicit the coupled roles of hardware and software design. Second, we organize the field around five coupled design questions: (i) what neural representations best capture speech intention across overt, mimed, and imagined speech; (ii) where and how neural activity should be recorded to balance coverage, resolution, chronic stability, and surgical feasibility; (iii) how datasets and annotations in experiments should be designed to support cross-subject transfer, multilingual generalization, and reproducible benchmarking; (iv) which decoding architectures and intermediate representations best trade off accuracy, latency, robustness, and interpretability; and (v) how systems should be evaluated, governed, and translated while preserving user agency, identity, and safety. Third, we synthesize concrete translational guidance, including a unified evaluation perspective and practical design priorities for user-centered clinical deployment. Consistent with these contributions, the following sections progress from a neural-mechanism framework to recording modality and experiment trade-offs, algorithmic advances, unified evaluation standards, and translational and ethical design pathways, followed by future directions and a synthesis of key priorities. More importantly, we use this synthesis to call for the field’s next phase: larger multilingual, multi-center datasets; longitudinal and standardized benchmarks; adaptive yet interpretable decoders; and prospective user-centered clinical validation with transparent reporting and data sharing.
@@ 2 Neural Mechanism Framework for Intracranial Language Decoding
As summarized in Fig. 3, this section links four neural-mechanistic themes to a common mechanism–recording–model–task framework, highlighting how biological constraints shape measurable signals and, in turn, decoder architecture choices.
@@ 2.1 Somatotopic Organization and Articulatory Kinematics in Sensorimotor Cortex
The sensorimotor cortex, particularly the ventral sensorimotor cortex (vSMC) and middle precentral gyrus (midPrCG), exhibits a spatially intermixed yet somatotopically organized tuning to vocal tract articulators (e.g., lips, tongue, jaw, larynx) [92, 125, 55]. At the single-electrode or single-neuron level, tuning to multiple articulators is intermixed, as observed in ventral premotor cortex (Brodmann area 6), where a small region ($3.2\times 3.2mm^{2}$) contains a rich, intermixed representation of speech articulators. This intermixed tuning is not limited to surface areas; even the dorsal ’hand knob’ area of the precentral gyrus shows broad tuning across phonemes, with many electrodes modulated for multiple phonemes, indicating spatially intermixed neural activity related to speech production [129]. Similarly, single neurons in the anterior precentral gyrus can be tuned to movements of multiple vocal-tract articulators or even whole-body movements [68, 64], and electrodes in vSMC encode coordinated trajectories involving multiple articulators rather than single articulator movements. Despite this intermixing, population-level somatotopic organization is evident: vSMC and midPrCG are somatotopically organized along the ventral–dorsal axis, representing different muscle groups controlling speech articulators [109], with an overall somatotopic dorsoventral arrangement of articulator representations (e.g., lips, jaw, tongue, larynx) on ventral pre- and post-central gyri, such as lip representation localized dorsally and tongue representation more broadly distributed ventrally, and clustering into four main groups corresponding to coronal, labial, dorsal, and vocalic articulatory configurations. The somatotopic organization is further supported by findings that during passive listening to speech sounds, differential activation occurs in the precentral motor cortex mirroring the somatotopic representation of the articulators that produce those phonemes [101], and that the spatial patterns of gamma band activity over sensorimotor cortex show distinct organizations for consonants versus vowels, confirming coordination of distinct articulatory representations [15]. The spatial arrangement of vocal tract articulators also shows somatotopic organization and intermixing, with linear transformations confirming dominant contributions of articulators like the tongue tip influenced by the tongue blade and dorsum [24], and somatotopically organized and spatially intermixed neural representations of orofacial articulatory kinematics observable through muscle activation patterns [39].
@@
Emergent population-level representations and harmonic oscillator dynamics are key features of speech encoding in the sensorimotor cortex. Neural population activity manifests low-dimensional kinematic state–space trajectories characterized by principal components that capture articulatory dynamics, showing biphasic vowel-consonant oscillatory trajectories consistent with harmonic oscillator dynamics [5]. Individual electrodes in vSMC encode diverse, functionally stereotyped articulatory kinematic trajectories (AKTs) with out-and-back trajectories exhibiting damped harmonic oscillator dynamics, corresponding to vocal tract constrictions defining phonemes [16]. Principal component analysis and clustering reveal cortical state-space organized into clusters corresponding to major oral articulators and vowels, with consonants and vowels occupying distinct regions consistent across subjects [9], and this study also revealed harmonic oscillator-like dynamics underlying the temporal progression of articulatory representations. Population-level neural dynamics during speech include a large condition-invariant signal (CIS) at movement initiation followed by rotatory (oscillatory) dynamics in the neural state during articulation [113], and the Directions Into Velocities of Articulators (DIVA) model identifies harmonic oscillator-like, population-level circuit dynamics emerging from interacting cortical areas during speech motor control [37]. These population-level representations enable high decoding accuracy: using naive Bayes classifiers on 1 s of neural population activity, 33 orofacial movements can be decoded with 92% accuracy, 39 phonemes with 62% accuracy, and 50 words with 94% accuracy [128], and decoding performance improves with the number of electrodes in a log-linear relationship.
@@
Controversies persist regarding the consistency of somatotopic patterns across individuals. Intracranial studies provide evidence for emergent population-level representations, but there is individual variability in spatial patterns. Speaker-dependent variations in sensorimotor cortex responses captured by SEEG are attributed mostly to electrode placement differences [6], and interindividual variability in neural activations and electrode placement contributes to inconsistencies in spatial activation patterns [86]. Variability is also observed in the extent and consistency of population dynamics and neural modulation between participants, with some showing stronger and more stable modulation than others. Additionally, while population-level articulatory dynamics are consistent across speakers [5], somatotopic articulatory representations in motor cortex are not consistently active or spatially organized during speech perception across individuals [22], and some consistencies in somatotopic tuning across subjects are observed alongside ongoing variability and inconsistencies. The generalizability of multi-subject models is supported by accommodating inter-subject variability in electrode placement, though performance variability remains [18]. Variability across individuals in somatotopic patterns and tuning also contributes to controversies regarding the consistency of these representations [95].

### 16 Mar 2026 v2

@@ Abstract
Intracranial language brain-computer interfaces (BCIs) offer a promising route for restoring communication in individuals with severe motor and speech impairments, but clinical translation remains limited by fragmented and heterogeneous evidence, as well as unresolved design trade-offs across neuroscience, hardware, algorithms, validation methods, and clinical integration. This review synthesizes recent progress across four key domains in intracranial speech neuroprosthetics: i) the neural mechanisms underlying overt, mimed, and imagined speech; ii) decision-oriented hardware comparisons of surgically implanted recording modalities, including microelectrode array (MEA), electrocorticography (ECoG), and stereotactic electroencephalography (SEEG); iii) experimental strategies for achieving cross-subject and multilingual generalization; and iv) advances in neural decoding, including sequence models, attention-based architectures (e.g., transformers), articulatory intermediate representations, and language-prior-assisted frameworks. We highlight persistent bottlenecks, including weak cross-subject transfer, long-term non-stationarity and recalibration burden, heterogeneous and non-comparable evaluation practices, limited naturalistic expressivity (especially for tonal/logosyllabic languages), and the low signal-to-noise ratio (SNR) of neural activity in covert speech decoding. Our contributions are threefold: (1) an end-to-end, decision-oriented synthesis that links neural representations to recording choices, experimental design, decoding model architectures, and translational constraints; (2) a structured framework organized around five coupled design questions, accompanied by a unified evaluation framework and a cross-linguistic, cross-task benchmark template that integrates objective, perceptual, expressive, conversational, and longitudinal metrics; and (3) user-centered translational guidance that includes agency-preserving shared control, verifiable performance priorities, and scenario-specific minimum viable system (MVP) profiles for differentiating between reliability-first home communication and fidelity-first conversational speech restoration. We conclude with a call for larger multilingual, multi-center longitudinal datasets; harmonized benchmarks; adaptive yet interpretable decoders; prospective clinical validation; and transparent data-sharing and reporting practices with robust ethical safeguards. These efforts are essential to accelerate the safe and equitable deployment of speech neuroprostheses.
@@ 1 Introduction
Speech is the most natural and efficient channel for human communication, yet it is among the first abilities compromised in a range of devastating neurological conditions [130, 18, 87]. Individuals with locked-in syndrome (LIS), advanced amyotrophic lateral sclerosis (ALS), brainstem stroke, or severe paralysis may remain cognitively intact while unable to produce intelligible speech or purposeful movement [99, 120, 87, 24]. For these patients, language BCIs, systems that decode intended linguistic content (and ideally expressive prosody) directly from neural activity, offer a pathway to restore communication with greater speed, privacy, and user agency than conventional assistive technologies [130, 81, 111, 82, 13].
Over the past decade, intracranial neural recordings have catalyzed rapid progress in language BCI research. Compared with non-invasive modalities [28, 142, 79, 61, 75, 46], intracranial approaches provide higher SNR and access to high-frequency activity that tightly tracks local cortical computations [45, 12], enabling decoding at the levels of phonemes, words, and continuous speech [72, 71, 130]. In parallel, the field has shifted from proof-of-concept demonstrations toward system-level designs that must operate reliably in real-world settings: across days and months, across electrode configurations, across languages and speaking styles, and across the spectrum from overt to silently mouthed or imagined speech [80, 4]. As summarized in the lane-based timeline in Fig. 1, this evolution represents not only a chronology of decoder improvements but a broader shift toward chronic, online systems and frameworks that explicitly address cross-subject transfer, electrode heterogeneity, and deployment constraints. These requirements are more demanding than those of laboratory demonstrations and necessitate the tight integration of neuroscience, hardware, machine learning, evaluation methodology, and clinical translation.
@@
On the technology side, intracranial language BCIs sit at the intersection of multiple recording modalities, each with fundamentally different trade-offs. Intracortical MEAs can capture high-dimensional spiking activity with fine spatial granularity but typically cover limited cortical territory and face chronic stability challenges [130, 13]. ECoG grids and depth electrodes such as SEEG offer broader coverage and clinically established implantation pathways, yet they sample neural activity at different spatial scales and with different noise characteristics [72, 71, 20]. The “best” modality therefore depends on the intended use case (research vs. home use), surgical feasibility, target neural features, and acceptable calibration burden. Experiment design is another critical consideration. The field has evolved from small, single-subject datasets with limited linguistic diversity toward larger, multi-subject datasets that enable cross-subject learning and benchmarking. However, challenges remain in standardizing data collection protocols, ensuring multilingual coverage, and producing rich annotations that capture not only lexical content but also prosody, emphasis, and affective features [123, 101, 35, 74, 124, 60].
@@
Algorithmically, language BCIs have evolved from feature-engineered pipelines and frame-wise classifiers to Seq2Seq models that jointly learn neural feature extraction, alignment, and decoding [82]. Advances such as transformers, convolutional temporal models, and hybrid architectures have enabled larger vocabularies, streaming or online decoding, and improved utilization of long-range context. Additionally, biologically grounded articulatory intermediates provide structural constraints that enhance data efficiency and generalization [20, 19]. More recently, language models and other linguistic priors have been integrated to reduce error rates and increase output coherence, raising new questions about controllability, bias, latency, and attribution of “who said what” in shared-control system [130, 13]. These trends underscore the importance of evaluating models not only in terms of accuracy but also considering calibration requirements, real-time constraints, robustness to neural non-stationarity, and interpretability.
@@
Despite these advances, several core bottlenecks continue to hinder clinical translation of intracranial language BCIs. First, poor cross-subject generalization persists: many decoders remain subject-specific, do not transfer across electrode layouts, and cannot fully leverage data across participants or centers [20, 19, 82]. Second, long-term stability of both neural signals and end-to-end performance remains a central challenge, with MEAs often requiring frequent recalibration and ECoG/SEEG systems still necessitating rigorous chronic validation under naturalistic use [80, 71, 72, 4, 13]. Third, evaluation practices remain heterogeneous: inconsistent use of word error rate (WER), phoneme error rate (PER), latency, information-transfer measures, and human perceptual scoring makes cross-task comparisons difficult and can obscure clinically meaningful differences [31, 133, 21, 116]. Fourth, limited naturalistic expressivity restricts many systems to lexical content, while paralinguistic features such as prosody, intonation, emphasis, and affect are less consistently modeled. This challenge is amplified for tonal languages (e.g., Mandarin, Cantonese) where pitch contours and syllable structure are integral to lexical meaning [5, 101, 35, 74, 124, 60, 11, 22]. Finally, covert speech decoding (silent articulation or imagined speech) poses unique difficulties, as neural signals are typically weaker, more variable, and less time-locked to measurable acoustic ground truth [76, 52].
@@
Despite significant recent progress, the field of intracranial language BCI remains highly fragmented across neuroscience, hardware development, algorithmic design, evaluation methods, and clinical translation. This fragmentation makes it difficult to compare studies, identify clinically meaningful trade-offs, and derive actionable design principles. In particular, progress towards clinical translation is hindered by several unresolved challenges, such as poor cross-subject generalization, limited evidence of long-term stability, heterogeneous evaluation metrics, and insufficient consideration of naturalistic expressivity and covert speech. To address this gap, this review makes three core contributions. First, we provide an end-to-end, decision-oriented synthesis that links neural mechanisms, recording methodologies, experiment design, decoding models, evaluation strategies, and translational deployment into a cohesive framework. Fig. 2 previews this end-to-end pipeline, tracing the path from neural recordings (e.g., ECoG, EEG, MEA) through feature extraction and neural decoding to outputs including phonemes, text, speech features, and synthesized speech, while making explicit how hardware and software choices jointly shape each stage. Second, we organize the field around five interconnected design questions that span the entire BCI development pathway: (i) What neural representations best capture speech intention across overt, mimed, and imagined speech? (ii) Where and how should neural activity be recorded to balance coverage, resolution, chronic stability, and surgical feasibility? (iii) How should experimental datasets and annotations be designed to enable cross-subject generalization, multilingual applicability, and reproducible benchmarking? (iv) Which decoding architectures and intermediate representations best trade off accuracy, latency, robustness, and interpretability? and (v) How should systems be evaluated, governed, and translated while preserving user agency, identity, and safety? Third, we synthesize concrete translational guidance, including a unified evaluation framework and practical design priorities for user-centered clinical deployment. Consistent with these contributions, the following sections progress from a foundational framework of neural mechanisms to the trade-offs inherent in recording modalities and experimental design, followed by algorithmic advances, unified evaluation standards, and pathways for translational and ethical development. We then conclude with future directions and a synthesis of key priorities. More importantly, we use this synthesis to call for the field’s next phase: larger multilingual and multi-center datasets; longitudinal and standardized benchmarks; adaptive yet interpretable decoders; and prospective user-centered clinical validation with transparent reporting and data sharing.
@@ 2 Neural Mechanism Framework for Intracranial Language Decoding
As summarized in Fig. 3, this section links four neural-mechanistic themes to a common mechanism–recording–model–task framework, highlighting how biological constraints shape measurable signals and, in turn, influence choices in decoder architecture.
@@ 2.1 Somatotopic Organization and Articulatory Kinematics in Sensorimotor Cortex
Within the sensorimotor cortex, the ventral sensorimotor cortex (vSMC) and middle precentral gyrus (midPrCG) exhibit spatially intermixed but somatotopically organized tuning to vocal tract articulators, including the lips, tongue, jaw, and larynx [94, 127, 57]. This intermixing extends to the single-electrode and single-neuron level: in the ventral premotor cortex (Brodmann area 6), even a tiny region of just $3.2\times 3.2mm^{2}$ exhibits a dense and overlapping representation of multiple speech articulators. This pattern is not confined to ventral speech areas: even the dorsal ’hand knob’ of the precentral gyrus shows broad tuning across phonemes, with many electrodes modulated for multiple phonemes, further evidence that neural activity related to speech production is spatially intermixed [131]. Similarly, single neurons in the anterior precentral gyrus can be tuned to movements of multiple vocal-tract articulators or even whole-body movements [70, 66], and electrodes in vSMC encode coordinated trajectories involving multiple articulators rather than single articulator movements. Despite this intermixing, population-level somatotopic organization is evident. The vSMC and midPrCG are somatotopically organized along the ventral–dorsal axis, representing different muscle groups that control the speech articulators [111]. This results in an overall dorsoventral arrangement of articulator representations (e.g., lips, jaw, tongue, larynx) on the ventral pre- and post-central gyri. For example, the lips are represented more dorsally, while the tongue is distributed more broadly across ventral areas. These representations cluster into four main groups, corresponding to coronal, labial, dorsal, and vocalic articulatory configurations. The somatotopic organization is further supported by two lines of evidence. First, during passive listening to speech sounds, differential activation occurs in the precentral motor cortex that mirrors the somatotopic representation of the articulators used to produce those phonemes [103]. Second, spatial patterns of gamma band activity over the sensorimotor cortex show distinct organizations for consonants versus vowels, confirming the coordination of distinct articulatory representations [16]. The spatial arrangement of vocal tract articulators shows both somatotopic organization and intermixing, with linear transformations confirming the dominant contributions of articulators like the tongue tip, which is influenced by the tongue blade and dorsum [25], and with somatotopically organized yet spatially intermixed neural representations of orofacial articulatory kinematics observable through muscle activation patterns [40].
@@
Emergent population-level representations and harmonic oscillator dynamics are key features of speech encoding in the sensorimotor cortex. Neural population activity yields low-dimensional kinematic state–space trajectories, with principal components capturing articulatory dynamics. These trajectories exhibit biphasic vowel-consonant oscillatory trajectories consistent with harmonic oscillator dynamics [5]. Individual electrodes in vSMC encode diverse, functionally stereotyped articulatory kinematic trajectories (AKTs). These out-and-back trajectories exhibit damped harmonic oscillator dynamics and correspond to the vocal tract constrictions that define phonemes [17]. Principal component analysis and clustering reveal that the cortical state-space is organized into clusters corresponding to major oral articulators and vowels, with consonants and vowels occupying distinct regions consistent across subjects [9]. This study also revealed harmonic oscillator-like dynamics underlying the temporal progression of articulatory representations. Population-level neural dynamics during speech include a large condition-invariant signal (CIS) at movement initiation, followed by rotatory (oscillatory) dynamics in the neural state during articulation [115]. The Directions Into Velocities of Articulators (DIVA) model identifies harmonic oscillator-like, population-level circuit dynamics emerging from interacting cortical areas during speech motor control [38]. These population-level representations enable high decoding accuracy: using naive Bayes classifiers on 1 s of neural population activity, 33 orofacial movements can be decoded with 92% accuracy, 39 phonemes with 62% accuracy, and 50 words with 94% accuracy [130]. Decoding performance improves with the number of electrodes in a log-linear relationship.
@@
Controversies persist regarding the consistency of somatotopic patterns across individuals. Intracranial studies provide evidence for emergent population-level representations, but there is individual variability in spatial patterns. Speaker-dependent variations in sensorimotor cortex responses captured by SEEG are attributed mostly to differences in electrode placement [6], with interindividual variability in neural activations and electrode placement contributing to inconsistencies in spatial activation patterns [88]. Variability is also observed in the extent and consistency of population dynamics and neural modulation between participants, with some showing stronger and more stable modulation than others. Additionally, while population-level articulatory dynamics are consistent across speakers [5], somatotopic articulatory representations in motor cortex are not consistently active or spatially organized during speech perception across individuals [23], and some consistencies in somatotopic tuning across subjects are observed alongside ongoing variability and inconsistencies. The generalizability of multi-subject models is supported by accommodating inter-subject variability in electrode placement, though performance variability remains [19]. Variability across individuals in somatotopic patterns and tuning also contributes to controversies regarding the consistency of these representations [97].